OpenStudy (samsan9):

what are the coordinates of the point of intersection of the two lines on the graph? of {10x+4y=24 {-6x-4y=8

OpenStudy (anonymous):

use elimination method 10 x + 4y = 24 -6x - 4y = 8 _________________ 4x = 32 x= 8 Now plug x into 10x + 4y = 24. find y

OpenStudy (samsan9):

i got the coordinated (8,-14) but idk how to graph this

OpenStudy (anonymous):

to graph it? Do you know how to find x and y intecepts? That's the easiest way to graph

OpenStudy (samsan9):

sorry i was eating and i got y=-10x/4+6 can you check if this is true?

OpenStudy (anonymous):

Yes :), though I suggest knowing how to graph x,y intercepts. It'd take you like 2 minutes to graph this system with that.

OpenStudy (samsan9):

can you show me the quicker and more efficient way?

OpenStudy (anonymous):

Sure, let's work with 10x + 4y = 24 find x intercept: plug in 0 = y. What do you get?

OpenStudy (anonymous):

10x + 0 = 24 x= 24/10 x= 2.4

OpenStudy (anonymous):

so there you go! You have point (2.4 , 0 ) on the first graph. Now let's make that x = 0 10(0) + 4y = 24

OpenStudy (anonymous):

tell me what you get for y.

OpenStudy (samsan9):

y=6

OpenStudy (anonymous):

yes, now you have ( 2.4 , 0 ) and ( 0 , 6). Plot those two points? then draw a straight line across it.

OpenStudy (anonymous):

across those points. That's how you graph the first equation :) Moving on to the second equation.

OpenStudy (anonymous):

-6x -4y = 8 find x intercept: let y = 0 -6x -4(0) = 8 what do you get for x

OpenStudy (anonymous):

Just work these out, then I'll prove that this works :)

OpenStudy (anonymous):

x= -8/6 x= -1.33 So you have (-1.33, 0 ) now let x = 0 -6(0) -4y = 8 y= -2 so you have (0,-2) plot these point for the 2nd equation. Make a straight line across these 2 points. Voila.

OpenStudy (samsan9):

y=-2 x=4/3

OpenStudy (samsan9):

sorry i forgot that we don't simplify 8/-6

OpenStudy (anonymous):

basically, plugging in 0 for x and y is the easiest way to find the answers :)

OpenStudy (anonymous):

you don't have to worry about both variables or extra numbers

OpenStudy (samsan9):

oh ok well i would have to graph this tomorrow since i have teacher scored trest but i have the same questions and this was one of them

OpenStudy (anonymous):

Ok, so just remember the points? haha. Jk, do it again with this method will save you time :)